OpenStudy (anonymous):

Solve for p: 3q= 5/6s(3z/t+5m/p)

OpenStudy (campbell_st):

so is the problem \[3q = \frac{5}{6s}(\frac{3z}{t} + \frac{5m}{p})\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

The first step is distributive property but everything is confusing after

OpenStudy (agent0smith):

Don't use distributive property first.

OpenStudy (anonymous):

Okay thanks So what's first?

OpenStudy (campbell_st):

ok... this is a long question.... mutliply both sides of the equation by 6s so \[18qs = 5(\frac{3z}{t} + \frac{5m}{p})\] now divide both sides by 5 \[\frac{18qs}{5} = \frac{3z}{t} + \frac{5m}{p}\] subtract 3z/t from both sides of the equation \[\frac{18qs}{5} - \frac{3z}{t} = \frac{5m}{p}\] hope this makes sense so far

OpenStudy (anonymous):

1 sec pls

OpenStudy (campbell_st):

the next step I'd recommend is put everything on the left hand side over a common denominator... just makes it easier...

OpenStudy (anonymous):

Then cancel out?

OpenStudy (campbell_st):

well start with the common denominator it will be 5t \[\frac{18qs}{5} \times \frac{t}{t} - \frac{3z}{t} \times \frac{5}{5} = \frac{5m}{p}\] so you get \[\frac{18 qst - 15z}{5t} = \frac{5m}{p}\] does that make sense..?

OpenStudy (anonymous):

Where does t/t come from?

OpenStudy (campbell_st):

well its equal to 1... and you are using it to get the common denominator of 5t

OpenStudy (campbell_st):

happy to go on..?

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

Ok I see

OpenStudy (campbell_st):

ok... here is a short cut... take the reciprocal of both fractions... \[\frac{5t}{(18qst - 15z)} = \frac{p}{5m}\] does that make sense...?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

It can't be reduced?

OpenStudy (campbell_st):

last step... multiply both sides of the equation by 5m \[\frac{5t \times 5m}{(18qst - 15z)} = p\] I'll let you finish it off... please go back and check the steps.

OpenStudy (anonymous):

Can 5t be cancelled out?

OpenStudy (anonymous):

25mt

OpenStudy (campbell_st):

nope... because 5t needs to be a factor of both terms in the denominator the only common factor in the denominator is 3 the 2nd part of the denominator, 15z, doesn't comtain a t and 18qst doesn't have 5 as a factor...

OpenStudy (campbell_st):

so the final answer is \[p = \frac{25mt}{(18qst - 15z)}\]

OpenStudy (anonymous):

I just found the answer Distributive property would've been easier then multiplied the COmmon demon. by 3q 18pstq=15pz+25mt P=25mt/18stq-15z

OpenStudy (anonymous):

Same answer but shorter

OpenStudy (campbell_st):

well its the same answer... and if you know fractions then the 1st step could have been divide both sides by 5/6s.... which is the same as multiplying by the reciprocal...

OpenStudy (anonymous):

Either way, wouldn't it be easier using distribute property?

OpenStudy (anonymous):

Shorter and simple

OpenStudy (campbell_st):

hope it all makes sense... good luck

OpenStudy (agent0smith):

ehhh... distributive property seems like more work, since you first have to distribute the 5/6s... it's already factored out, and distributing is undoing the work that was done to factor it out in the first place.

OpenStudy (anonymous):

I appreciate both of your help but i understand it so much better when I distribute Although I never understood the problem until it was explained, thank you

OpenStudy (agent0smith):

Fair enough, you can do it whichever way you prefer, distributing is still a valid way to re-arrange it.